# EqNIO: Subequivariant Neural Inertial Odometry

### v1

@@ 4.3 Basic Layers
Solving the above equation amounts to finding the eigenspace of the left-most matrix with eigenvalue 1. Such analysis for $R_{2\times 2}\in SO(2)$ yields $W_{SO(2)}=w_{1}I_{2\times 2}+w_{2}H_{2\times 2}$, where $H_{2\times 2}$ denotes a $90$ degree counter-clockwise rotation in 2D, and $w_{1},w_{2}\in\mathbb{R}$ are learnable weights. Similarly, for $O(2)$ we find $W_{O(2)}=w_{1}I_{2\times 2}$. Vectorizing this linear mapping to multiple input and output vector features we have the following $SO(2)$ and $O(2)$ equivariant linear layers:
$SO(2)$ $:\quad v^{\text{out}}=v^{\text{in}}W_{1}+Hv^{\text{in}}W_{2}$ (4)
@@
Previous works [47, 46] propose various nonlinearities such as norm-nonlinearity, tensor-product nonlinearity, and gated nonlinearity for $SO(3)$ and $O(2)$ equivariance in an equivariant convolutional way; while Deng et al. [20] applies per-point nonlinearity for vector features only. Since we apply convolutions over time, and there is no transformation on this dimension, we don’t apply nonlinearity in a convolutional way but apply a nonlinearity pointwise. On the other hand, since our data includes both scalars and vectors unlike [20], we want to mix these two types of features. Therefore, we adapt gated nonlinearity [47] to pointwise nonlinearity. Specifically, for $n$ vector and scalar features $v^{\text{in}}\in\mathbb{R}^{n\times 2\times C},s^{\text{in}}\in\mathbb{R}^{n\times C}$, we concatenate the vector norm features $\|v^{\text{in}}\|\in\mathbb{R}^{n\times C}$ with the scalar features $s^{\text{in}}$, and run a single MLP yielding an output of size $2C$. We then split this output into new norm features $\gamma\in\mathbb{R}^{n\times C}$ and new activations $\beta\in\mathbb{R}^{n\times C}$ which we modulate with a non-linearity $s^{\text{out}}=\sigma(\beta)$. Finally, we rescale the original vector features according to the new norm. In summary, the update looks like
@@ 4.4 Equivariant Covariance
Our loss function follows the previous work [33] as $\mathcal{L}_{MSE}(\hat{d},d)=\|d-\hat{d}\|_{2}$ in first stage, and $\mathcal{L}_{MLE}(\hat{d},d)=(d-\hat{d})^{T}\Sigma^{-1}(d-\hat{d})$ in the second stage when $\mathcal{L}_{MSE}$ converges. TLIO [33] opted for a diagonal covariance matrix instead of parameterizing the off-diagonal elements using Pearson correlation, to enhance training stability [36]. However, such an assumption breaks the equivariance of the covariance and, therefore, breaks the invariance of $\mathcal{L}_{MLE}$. This issue arises because a rotation of the input invariably alters the covariance, preventing it from remaining diagonal consistently.
@@ 5 Experiments
To demonstrate the effectiveness of canonicalizing the IMU data using the equivariant framework, we apply it to two types of neural inertial navigation systems: 1) end-to-end deep learning approach (RoNIN), and 2) filter-based approach with a learned prior (TLIO). We use the specific test datasets only to provide a fair comparison with prior work. We do extensive ablations in Section 6 to illustrate the usefulness of modeling symmetry and validate our design choices for the subequivariant framework.
The equivariant framework is applied to TLIO and RONIN architectures. For the TLIO architecture, we train on TLIO Dataset [33] and test on TLIO Dataset and Aria Everyday Activities (Aria) Dataset  [34]. For the RONIN architecture, we test on the three popular pedestrian datasets RONIN [30], RIDI [52] and OxIOD [11]. (For more details on the datasets refer to AppendixA.2. The details of the equivariant network are provided in AppendixA.3 and visualized in Figure 2(b).) TLIO estimates orientation, position, velocity, IMU biases, and noise using an EKF which takes displacement and uncertainty predicted by a neural network as measurement update. RONIN regresses only the 2D velocity. Both architectures take IMU samples in a gravity-aligned frame without gravity compensation as input to the neural network. The Neural Network performance (indicated with *) is evaluated using three metrics Mean Squared Error (MSE) in $m^{2}$, Absolute Translation Error (ATE) in $m$, and Relative Translation Error (RTE) in $m$, and EKF results are evaluated using ATE in $m$, RTE in $m$ and Absolute Yaw Error (AYE) in degrees. Details of these metrics are provided in the Appendix A.5. Eq F. refers to EqNIO, i.e., the application of Equivariant Framework.
@@ 5.1 TLIO Architecture
[…]
Table 1: Neural network results (no EKF) are labeled with *. MSE are multiplied by 100. TLIO-N is TLIO trained without yaw augmentations. $\dagger$ represents TLIO+.
Figure 5: EKF Results for TLIO on TLIO and Aria Datasets visualized with a box plot.
Figure 6: Visualization trajectory for TLIO. $1.a$ and $1.b$ are easy trajectories; $2.a$ and $2.b$ are mid-level hard trajectories; $3.a$ and $3.b$ are unusual motions not present in the training set are performed.
In Table 1, we compare baseline TLIO, trained with yaw augmentations as mentioned in their paper, TLIO without the augmentation, and our two methods applied to TLIO. As seen in Table  1, the Eq F. $O(2)$ has better performance on all three metrics by a huge margin with the baseline TLIO model, and the Eq F. $SO(2)$ model closely follows. The performance of our methods is consistent across TLIO and Aria Datasets illustrating our generalization ability. Table  1 and Figure  5 shows our method surpasses the baseline on most metrics while remaining comparable in AYE. The superior performance of our model when combined with EKF is attributed to its generalization ability when the orientation estimate is not very accurate as well as the the equivariant covariance predicted by the network. The performance of our method on AYE is comparable to TLIO with a difference of magnitude 0.02 degrees and the equivariant network is more tolerant to yaw perturbations. The trajectory plots for trajectories from the TLIO test dataset are as seen in Figure 6 and Appendix A.6.
@@ 5.2 RONIN Architecture
RONIN + Eq F. $SO(2)$ 5.18 4.35 3.67 2.72 0.86 1.59 0.63 1.39 1.22 2.39
RONIN + Eq F. $O(2)$ 4.42 3.95 3.32 2.66 0.82 1.52 0.70 1.41 1.28 2.10
Table 2: Results for RONIN. We follow the original paper to report results with two decimal places. RONIN-U, RONIN-S, RIDI-T, RIDI-C, OXOID refers to RONIN Unseen Dataset, RONIN Seen Dataset, RIDI Test Dataset, RIDI Cross Subject Dataset and OxIOD Dataset respectively.
Table 2 compares the RONIN [30], RIO [8], and our methods. RONIN and RIO results follow their respective publications. We trained RONIN using only the available 50% dataset. RIO identifies the original RONIN model trained on this dataset as B-ResNet, and J-ResNet as RONIN optimized for both MSE loss on velocity prediction and cosine similarity with an equivariance constraint modeled using an auxiliary loss. RIO also adapted these models using their Adaptive Test-Time-Training strategy, reported as B-ResNet-TTT and J-ResNet-TTT. They do not provide results for RONIN Seen Dataset (RONIN-S) or RIDI Cross Subject Dataset (RIDI-C). Our equivariant methods significantly outperform the original models, with the Eq F. $O(2)$ model converging at 38 epochs compared to over 100 in RONIN. This demonstrates the superior generalization of our fully equivariant architecture. RIO’s approach, involving multiple data rotations, test optimization, and a deep ensemble at test time, results in higher computational and memory costs.
@@ 6 Ablation Study
[…]
Table 3: Ablation Studies. TLIO-NQ is TLIO with non-equivariant frames. TLIO-PCA is TLIO using PCA to predict handcrafted frames. TLIO-N is TLIO trained without yaw augmentations. Eq F. $\bm{SO(2)}$+S is Equivaraiant frame with circular covariance. Eq F. $\bm{SO(2)}$+P is $SO(2)$ equivariant networks with pearson parameterized covariance. Eq CNN is fully equivariant CNN. The same naming conventions apply to the $O(2)$ variants. $\dagger$ represents TLIO+.
In this section, we investigate and motivate the necessity for incorporating equivariance in inertial odometry, the choice of equivariant architecture and covariance. We present all the ablation on TLIO in Table 3 for the neural network and the EKF results. Appendix A.7 contains the results of evaluating all the above models separately on a test dataset augmented with rotations and/or reflections. Appendix A.10 presents the ablation on IMU input sequence length and lastly, in Appendix A.11 we present sensitivity analysis to gravity direction using 5 discrete angles.
Baseline Ablation: Is yaw augmentation needed when the input is in a local gravity-aligned frame? We trained TLIO both with and without yaw augmentation using identical hyperparameters and the results in Table 3 revealed that augmentation enhances the network’s generalization, improving all metrics for the Aria dataset. This underscores the importance of equivariance for network generalization. Does a Deeper TLIO with a comparable number of parameters match the performance of equivariant methods? We enhanced the residual depth of the original TLIO architecture from [2, 2, 2, 2] to [3, 3, 3, 3] to align its number of parameters with our Eq F. $SO(2)$ model. Despite having fewer parameters due to the removal of the orthogonal basis in the $SO(2)$ vector neuron-based architecture, the Eq F. $O(2)$ model still outperformed the augmented TLIO. The data from Table 3 demonstrate that merely increasing the network’s size, without integrating true equivariance, is insufficient for achieving precise inertial odometry.
@@
Frame Ablation: Can a non-equivariant MLP predict meaningful frames? We trained TLIO with augmentation and identical hyperparameters alongside an additional MLP mirroring the architecture of our method to predict a frame. We observed that the TLIO-NQ model tends to overfit to the TLIO dataset, and the predicted frames were not meaningful, as illustrated in Figure 7. Can frames predicted using PCA (handcrafted equivariant frame) achieve the same performance? PCA frames underperform on the Aria dataset and perform worse than the original TLIO, likely due to PCA’s noise sensitivity and the production of non-smooth frames, as shown in Figure 7. Additionally, PCA cannot distinguish between $SO(2)$ and $O(2)$ transformations. Figure 7 also shows that $O(2)$ does not have frames as smooth as $SO(2)$ as the reflected bends have reflected frames.
@@ A.6 Visualization of TLIO results
Figure 8: Only Neural Network Results for TLIO on TLIO and Aria Datasets visualized with a box plot.
Figure 9: Only Neural Network Results for TLIO on TLIO and Aria Datasets visualized with a box plot.
Figure 10: EKF Results for TLIO on TLIO and Aria Datasets visualized with a box plot.
Figure 11: More Visualizations of predicted and ground-truth trajectories from Tlio Test Dataset
Figure 12: More Visualizations of predicted and ground-truth trajectories from Tlio Test Dataset

### v2

@@ 4.3 Basic Layers
Solving the above equation amounts to finding the eigenspace of the left-most matrix with eigenvalue 1. Such analysis for $R_{2\times 2}\in SO(2)$ yields $W_{SO(2)}=w_{1}I_{2\times 2}+w_{2}R_{90}$, where $R_{90}$ denotes a $90$ degree counter-clockwise rotation in 2D, and $w_{1},w_{2}\in\mathbb{R}$ are learnable weights. Similarly, for $O(2)$ we find $W_{O(2)}=w_{1}I_{2\times 2}$. Vectorizing this linear mapping to multiple input and output vector features we have the following $SO(2)$ and $O(2)$ equivariant linear layers:
$SO(2)$ $:\quad v^{\text{out}}=v^{\text{in}}W_{1}+R_{90}v^{\text{in}}W_{2}$ (4)
@@
Previous works [47, 46] propose various nonlinearities such as norm-nonlinearity, tensor-product nonlinearity, and gated nonlinearity for $SO(3)$ and $O(2)$ equivariance in an equivariant convolutional way; while Deng et al. [20] applies per-point nonlinearity for vector features only. Since we apply convolutions over time, and there is no transformation on this dimension, we don’t apply nonlinearity in a convolutional way but apply a nonlinearity pointwise. On the other hand, since our data includes both scalars and vectors unlike [20], we want to mix these two types of features. Therefore, we adapt gated nonlinearity [47] to pointwise nonlinearity. Specifically, for $n$ vector and scalar features $v^{\text{in}}\in\mathbb{R}^{n\times 2\times C},s^{\text{in}}\in\mathbb{R}^{n\times C}$, we concatenate the vector norm features $\|v^{\text{in}}\|\in\mathbb{R}^{n\times C}$ with the scalar features $s^{\text{in}}$, and run a single MLP yielding an output of size $n\times 2C$. We then split this output into new norm features $\gamma\in\mathbb{R}^{n\times C}$ and new activations $\beta\in\mathbb{R}^{n\times C}$ which we modulate with a non-linearity $s^{\text{out}}=\sigma(\beta)$. Finally, we rescale the original vector features according to the new norm. In summary, the update looks like
@@ 4.4 Equivariant Covariance
Our loss function follows the previous work [33] as $\mathcal{L}_{MSE}(\hat{d},d)=\|d-\hat{d}\|_{2}$ in the first stage, and $\mathcal{L}_{MLE}(\hat{d},d)=(d-\hat{d})^{T}\Sigma^{-1}(d-\hat{d})$ in the second stage when $\mathcal{L}_{MSE}$ converges. TLIO [33] opted for a diagonal covariance matrix instead of parameterizing the off-diagonal elements using Pearson correlation, to enhance training stability [36]. However, such an assumption breaks the equivariance of the covariance and, therefore, breaks the invariance of $\mathcal{L}_{MLE}$. This issue arises because a rotation of the input invariably alters the covariance, preventing it from remaining diagonal consistently.
@@ 5 Experiments
To demonstrate the effectiveness of canonicalizing the IMU data using the equivariant framework, we apply it to two types of neural inertial navigation systems: 1) end-to-end deep learning approach (RONIN), and 2) filter-based approach with a learned prior (TLIO). Both architectures take IMU samples in a gravity-aligned frame without gravity compensation as input to the neural network. While RONIN regresses only the 2D velocity, TLIO estimates orientation, position, velocity, IMU biases, and noise using an EKF which takes displacement and uncertainty predicted by a neural network as measurement update.
In this section, we train our framework applied to TLIO architecture on TLIO Dataset [33] and test on TLIO Dataset and Aria Everyday Activities (Aria) Dataset  [34]. Our method applied to RONIN is trained on RONIN [30]. RONIN, RIO’s methods, and our two frameworks are tested on the three popular pedestrian datasets RONIN [30], RIDI [52] and OxIOD [11]. We use the specific test datasets only to provide a fair comparison with prior work. See AppendixA.2 for more dataset details, and see AppendixA.3 and Figure 2(b) for more details and visualizations of the equivariant network. Further, in Section 6 we conduct extensive ablation illustrating the usefulness of modeling symmetry and validate our design choices for the subequivariant framework.
The Neural Network performance (indicated with *) is evaluated using three metrics: Mean Squared Error (MSE) in $m^{2}$, Absolute Translation Error (ATE) in $m$, and Relative Translation Error (RTE) in $m$. For evaluating only the NN, the trajectory is reconstructed via cumulative summation from the initial state as done in prior work. In case of TLIO architecture, we additionally evaluate the overall performance after integrating the NN with the EKF and report the ATE in $m$, RTE in $m$, and Absolute Yaw Error (AYE) in degrees. Details of these metrics are provided in the Appendix A.5. In the following text and tables, Eq F. will refer to EqNIO, i.e., the application of Equivariant Framework.
@@ 5.1 TLIO Architecture
TLIO + Eq F. $\mathbf{SO(2)}$ 3.194 1.480 2.401 0.490 0.501 2.428 2.457 1.178 1.864 0.449 0.484 2.084
TLIO + Eq F. $\mathbf{O(2)}$ 2.982 1.433 2.406 0.458 0.478 2.389 2.304 1.118 1.850 0.416 0.465 2.059
Table 1: Neural network results (no EKF) are labeled with *. MSE are multiplied by 100. TLIO-N is TLIO trained without yaw augmentations. A lower error indicates a better model. A lowest values are annotated with Red, while the second lowest with Orange. Our proposed methods are in bold.
Figure 5: The superior performance of our framework applied to TLIO architecture when compared to baseline TLIO trained with and without augmentations on TLIO and Aria Datasets visualized with a box plot. Blue, Orange, Green and Red indicate TLIO, TLIO-N, TLIO+Eq F. $SO(2)$ and TLIO+Eq F. $O(2)$.
Figure 6: Visualization of final estimated trajectories on TLIO Dataset by baseline TLIO (Blue), our best method applied to TLIO (TLIO+Eq F. $O(2)$)(Red), and the Ground-Truth trajectory (Black). $1.a$ and $1.b$ are easy trajectories; $2.a$ and $2.b$ are mid-level hard trajectories; $3.a$ and $3.b$ are unusual motions not present in the training set are performed.
In Table 1, we compare baseline TLIO, trained with yaw augmentations as mentioned in their paper (termed TLIO), TLIO without the augmentation (termed TLIO-N), and our two methods applied to TLIO (termed Eq.F. $SO(2)$ and $O(2)$). As seen in Table  1, Eq F. $O(2)$ has better performance on all three metrics comparing the NN (indicated by *) by a large margin of 56%, 12%, and 10% with the baseline TLIO model on MSE*, ATE*, and RTE* respectively, and the Eq F. $SO(2)$ model closely follows with 56%, 11% and 7% respectively on Aria Dataset. The performance of our methods is consistent across TLIO and Aria Datasets illustrating our generalization ability. Table  1 and Figure  5 shows our method surpasses the baseline on most metrics while remaining comparable in AYE. The superior performance of our model as compared to baseline TLIO when the NN is combined with EKF (i.e., performance on ATE, RTE and AYE metrics) is attributed to its generalization ability when the orientation estimate is not very accurate as well as the equivariant covariance predicted by the network. The AYE of our method is comparable to TLIO with a difference of magnitude 0.02 degrees and the equivariant network is more tolerant to yaw perturbations. The trajectory plots for trajectories from the TLIO test dataset are seen in Figure 6 and Appendix A.6.
@@ 5.2 RONIN Architecture
RONIN + Eq F. $\mathbf{SO(2)}$ 5.18 4.35 3.67 2.72 0.86 1.59 0.63 1.39 1.22 2.39
RONIN + Eq F. $\mathbf{O(2)}$ 4.42 3.95 3.32 2.66 0.82 1.52 0.70 1.41 1.28 2.10
Table 2: Results for RONIN. We follow the original paper to report results with two decimal places. RONIN-U, RONIN-S, RIDI-T, RIDI-C, OxIOD refers to RONIN Unseen Dataset, RONIN Seen Dataset, RIDI Test Dataset, RIDI Cross Subject Dataset and Oxford Inertial Odometry Dataset respectively. A lower error indicates a better model. The lowest values are annotated with Red, while the second lowest with Orange. Our proposed methods are in bold.
We compare our methods applied to RONIN with the original RONIN [30], 4 methods reported by RIO [8]- RIO B-ResNet is the original RONIN model retrained, RIO J-ResNet is optimized for both MSE loss on velocity prediction and cosine similarity with an equivariance constraint modeled using an auxiliary loss, RIO B-ResNet-TTT is original RONIN model adapted at test time using Adaptive Test-Time-Training strategy, and RIO J-ResNet-TTT is RIO J-ResNet adapted at test time using Adaptive Test-Time-Training strategy. Original RONIN is trained on 100% RONIN train dataset while RIO and our methods are trained on only the publicly available 50% of RONIN train dataset. RIO does not provide results for RONIN Seen Dataset (RONIN-S) or RIDI Cross Subject Dataset (RIDI-C). As Seen in Table 2, our equivariant methods significantly outperform the original RONIN by a large margin of 14% and 9% on ATE* and RTE* respectively even on the RONIN-U dataset. Our methods have better generalization as seen on RIDI-T and OxIOD, outperforming even RIO J-ResNet-TTT by a margin of 56% and 43% on ATE* and RTE* respectively on OxIOD Dataset. The Eq F. $O(2)$ model converges at 38 epochs compared to over 100 in RONIN implying faster network convergence with our framework as compared to data augmentation used in RONIN. This demonstrates the superior generalization of our strictly equivariant architecture. RIO’s approach, involving multiple data rotations, test optimization, and a deep ensemble at test time, would result in higher computational and memory costs as compared to our method.
@@ 6 Ablation Study
TLIO + Eq F. $SO(2)$+S 3.331 1.626 2.796 0.524 0.536 2.440 2.591 1.146 2.067 0.466 0.517 2.089
TLIO + Eq F. $SO(2)$+P 3.298 1.842 2.652 0.588 0.523 2.537 2.635 1.592 2.303 0.585 0.539 2.232
TLIO + Eq F. $\mathbf{SO(2)}$ 3.194 1.480 2.401 0.490 0.501 2.428 2.457 1.178 1.864 0.449 0.484 2.084
TLIO + Eq F. $O(2)$+S 3.061 1.484 2.474 0.462 0.481 2.390 2.421 1.175 1.804 0.421 0.458 2.043
TLIO + Eq F. $O(2)$+P 2.990 1.827 2.316 0.578 0.478 2.534 2.373 1.755 1.859 0.564 0.468 2.223
TLIO + Eq F. $\mathbf{O(2)}$ 2.982 1.433 2.406 0.458 0.478 2.389 2.304 1.118 1.849 0.416 0.465 2.059
Table 3: Ablation Studies. TLIO-NQ is TLIO with non-equivariant frames. TLIO-PCA is TLIO using PCA to predict handcrafted frames. TLIO-N is TLIO trained without yaw augmentations. Eq F. $\bm{SO(2)}$+S is Equivaraiant frame with circular covariance. Eq F. $\bm{SO(2)}$+P is $SO(2)$ equivariant networks with pearson parameterized covariance. Eq CNN is fully equivariant CNN. The same naming conventions apply to the $O(2)$ variants. A lower error indicates a better model. The lowest values are annotated with Red, followed by Orange and Yellow respectively. Our proposed methods are in bold.
In this section, we investigate and motivate the necessity for incorporating equivariance in inertial odometry, the choice of equivariant architecture and covariance. We present all the ablation on TLIO in Table 3 for the neural network and the overall performance when NN is integrated with the EKF. Appendix A.7 contains the results of evaluating all the above models separately on a test dataset augmented with rotations and/or reflections. Appendix A.10 presents the ablation on IMU input sequence length and lastly, in Appendix A.11 we present sensitivity analysis to gravity direction using 5 discrete angles.
Baseline Ablation: Is yaw augmentation needed when the input is in a local gravity-aligned frame? We trained TLIO both with and without yaw augmentation using identical hyperparameters and the results in Table 3 revealed that augmentation enhances the network’s generalization, improving all metrics for the Aria dataset with the lowest margin of 10% on AYE and highest margin of 65% for MSE*. This underscores the importance of equivariance for network generalization. Does a Deeper TLIO with a comparable number of parameters match the performance of equivariant methods? We enhanced the residual depth of the original TLIO architecture from 4 residual blocks of depth 2 each to 4 residual blocks with depth 3 each to align its number of parameters with our Eq F. $SO(2)$ model. Despite having fewer parameters due to the removal of the orthogonal basis in the $SO(2)$ vector neuron-based architecture, the Eq F. $O(2)$ model still outperformed the augmented TLIO. The data from Table 3 demonstrate that merely increasing the network’s size, without integrating true equivariance, is insufficient for achieving precise inertial odometry.
@@
Frame Ablation: Can a non-equivariant MLP predict meaningful frames? We trained TLIO with augmentation and identical hyperparameters alongside an additional MLP mirroring the architecture of our method to predict a frame and term this baseline TLIO-NQ. We observed that TLIO-NQ tends to overfit to the TLIO dataset, and the predicted frames were not meaningful, as illustrated in Figure 7. Can frames predicted using PCA (handcrafted equivariant frame) achieve the same performance? PCA frames underperform on the Aria dataset and perform worse than the original TLIO, likely due to PCA’s noise sensitivity and the production of non-smooth frames, as shown in Figure 7. Additionally, PCA cannot distinguish between $SO(2)$ and $O(2)$ transformations. Figure 7 also shows that $O(2)$ does not have frames as smooth as $SO(2)$ as the reflected bends have reflected frames.
@@ A.6 Visualization of TLIO results
Figure 8: The superior performance of our framework applied to TLIO architecture when compared to baseline TLIO trained with and without augmentations on TLIO and Aria Datasets visualized with a box plot. Blue, Orange, Green and Red indicate TLIO, TLIO-N, TLIO+Eq F. $SO(2)$ and TLIO+Eq F. $O(2)$. ATE, RTE and MSE indicate ATE*, RTE* and MSE* corresponding to only the NN results.
Figure 9: The superior performance of our framework applied to TLIO architecture when compared to baseline TLIO trained with and without augmentations on TLIO and Aria Datasets visualized with a box plot. Blue, Orange, Green and Red indicate TLIO, TLIO-N, TLIO+Eq F. $SO(2)$ and TLIO+Eq F. $O(2)$. ATE, RTE and MSE indicate ATE*, RTE* and MSE* corresponding to only the NN results. The whisker is extended to 1.5 * IQR (inter-quartile range).
Figure 10: The superior performance of our framework applied to TLIO architecture when compared to baseline TLIO trained with and without augmentations on TLIO and Aria Datasets visualized with a box plot. Blue, Orange, Green and Red indicate TLIO, TLIO-N, TLIO+Eq F. $SO(2)$ and TLIO+Eq F. $O(2)$. The whisker is extended to 1.5 * IQR (inter-quartile range).
Figure 11: More Visualizations of final estimated trajectories on TLIO Dataset by baseline TLIO (Blue), our best method applied to TLIO (TLIO+Eq F. $O(2)$)(Red), and the Ground-Truth trajectory (Black).
Figure 12: More Visualizations of final estimated trajectories on TLIO Dataset by baseline TLIO (Blue), our best method applied to TLIO (TLIO+Eq F. $O(2)$)(Red), and the Ground-Truth trajectory (Black).